OpenStudy (anonymous):

what is the length of the transverse axis of the conic section shown below ? (y+2)^2/16-(x-3)^2/9=1

OpenStudy (anonymous):

\[\large \frac{(y+2)^2}{16}-\frac{(x-3)^2}{9}=1\]

OpenStudy (anonymous):

What kind of hyperbola do you think it is?